OpenStudy (anonymous):

The boiling point of an aqueous solution is 102.42 °C. What is the freezing point? Constants can be found here.

OpenStudy (anonymous):

Since pure water boils at 100.00°C, the change in boiling point is ΔTb = (102.42 – 100.00)°C = 2.42 °C. The molal concentration, m (or b), and the van\'t Hoff factor, i, are constant, so solve the boiling-point-elevation formula for m·i. \[\Delta T _{b} = k _{b} \times m \times i\] \[mi = \frac{\Delta T _{b} }{ k _{b} }\] Then substitute that expression for m·i in the freezing-point-depression formula and solve for \[T _{f}\]